Mr. Pascual, an airline owner, decided to invest P2.5 million to fund his department handling  spare parts replacement. How long will it take his investment to accumulate to P3.325  million if the bank’s interest rate is 5.5%?

1
Expert's answer
2022-03-14T12:38:19-0400

The number of years it will take Mr Pascual to accumulate P3.325 at 5.5% interest rate is given as:

"N=\frac{log(\frac{FV}{PV})}{log(1+r)}"


"N=\frac{log(\frac{P3.325}{P2.5})}{log(1+0.055)}"

"N=5.32 \space{years}"